# Bilinear Causal Representation for Generalizable Offline Model-based Reinforcement Learning

### v1

@@ Abstract
Offline model-based reinforcement learning (MBRL) enhances data efficiency by utilizing pre-collected datasets to learn models and policies, especially in scenarios where exploration is costly or infeasible. Nevertheless, its performance often suffers from the objective mismatch between model and policy learning, resulting in inferior performance despite accurate model predictions. This paper first identifies the primary source of this mismatch comes from the underlying confounders present in offline data for MBRL. Subsequently, we introduce BilinEar CAUSal rEpresentation (BECAUSE), an algorithm to capture causal representation for both states and actions to reduce the influence of the distribution shift, thus mitigating the objective mismatch problem. Comprehensive evaluations on 18 tasks that vary in data quality and environment context demonstrate the superior performance of BECAUSE over existing offline RL algorithms. We show the generalizability and robustness of BECAUSE under fewer samples or larger numbers of confounders. Additionally, we offer theoretical analysis of BECAUSE to prove its error bound and sample efficiency when integrating causal representation into offline MBRL.
@@ 1 Introduction
Offline Reinforcement Learning (RL) has shown great promise in learning directly from historically collected datasets, especially in scenarios where active interaction is expensive or infeasible [1]. Specifically, offline model-based reinforcement learning (MBRL) [2, 3, 4], learning policies with an estimated world model, generally perform better than their model-free counterparts in long-horizon tasks such as self-driving vehicles [5], robotics [6], and healthcare [7]. However, offline RL suffers from distribution shift since the rollout data could sample from some unknown behavior policies that are sub-optimal or from slightly different environments compared to the deployment time [8].
@@
Although identifying distribution shift issues, many of the current offline MBRL works fail to model the shift in environment dynamics, which is ubiquitous and could cause catastrophic failure of trained policy at a slightly different deployment stage. Furthermore, since the learning objectives of the world models and policies are isolated from each other, a significant challenge in offline MBRL is objective mismatch [9, 10] problem (shown in Figure 1): models that achieve a lower training loss are not necessarily better for control performance. For example, in long-horizon planning tasks, the reward is sparse yet the prediction accuracy of the model may decay as the horizon enlarges and compounding error accumulates. Previous works [9, 10] have attempted to reduce such objective mismatch by jointly learning the model and policy. However, the performance is suboptimal in the absence of the underlying cause of objective mismatch [8].
In this work, we identify that the objective mismatch between model estimation and policy learning comes from two sources of distribution shift in offline MBRL: (1) shift between the online optimal policy and offline sub-optimal behavior policies, and (2) shift between the data collection environment and online testing environments. Unlike humans, who make decisions based on reasoning over task-relevant factors, models in offline RL memorize correlations without learning the causality. The sub-optimal behavior policies introduce spurious correlations [11] between actions and states, making the model memorize specific actions. When online testing environments differ from the data collection environment, the model could overfit spurious correlations in the state and fail to generalize to unseen states. Based on the analysis of the above mismatch, our work differs from previous work of causal model-based RL [12, 13] in that we model causality in both model and policy learning. We aim to avoid spurious correlation by discovering underlying structures between abstracted states and actions.
@@ 3.1 Causal Representation for ASC-MDP
The assumption 1 can also be interpreted as task independence in  [24], invariant state representation [25], or invariant action effect in [26].
@@ 4 Experiment Results
In this section, we conduct a comprehensive empirical evaluation of BECAUSE’s generalization performance in a diverse set of environments, covering different decision-making problems in the grid world, manipulation, and autonomous driving domains, shown in Appendix Figure 8.
@@ 4.1 Experiment Setting
We design 18 tasks in 3 representative RL environments. Agents need to acquire reasoning capabilities to receive higher rewards and achieve goals.
@@ 4.2 Experiment Results Analysis
  • RQ3: How does BECAUSE achieve the aforementioned generalizability by mitigating the objective mismatch problem in offline MBRL?
@@
Figure 6: The ablation studies between BECAUSE and its variants. We report the overall Success rate (%) over 9 in-distribution (I) and 9 out-of-distribution (O) tasks, respectively. Bold is the best.
@@
For RQ3, we aim to understand whether BECAUSE achieves higher performance by resolving the objective mismatch problem. We first collect two groups of trajectories: $\bm{\tau_{pos}}$ and $\bm{\tau_{neg}}$, each with positive reward (success) and negative reward (failure) in Unlock task with sparse goal-reaching reward. We want to have a model whose loss is informative for discriminating control results, that is, we wish ${\mathcal{L}}_{model}({\color[rgb]{0.44,0.72,0.93}\bm{\tau_{pos}}})<{\mathcal{L}}_{model}({\color[rgb]{0.96,0.49,0.43}\bm{\tau_{neg}}})$. According to our visualization in Figure 5, in Unlock-Expert and Unlock-Medium, the ratio of $\bm{\tau_{pos}}$ is much higher in BECAUSE than MOPO among the trajectories with low model loss. In Unlock-Random, the mismatch of the model and control objective is more significant, since the demonstration is poor in state coverage. MOPO cannot succeed even when the model loss is low, whereas our methods can. We perform a hypothesis test with $H_{0}:{\mathcal{L}}_{model}({\color[rgb]{0.44,0.72,0.93}\bm{\tau_{pos}}})<{\mathcal{L}}_{model}({\color[rgb]{0.96,0.49,0.43}\bm{\tau_{neg}}})$. In BECAUSE, this desired property is more significant than MOPO attributed to the causal representation we learn, indicating a reduction of objective mismatch. We attach detailed discussions for the mismatch evaluation in Appendix C.3 and Table 2.
4.3 Ablation Studies
We conducted ablation studies with three variants of BECAUSE and report the average success rate across nine in-distribution and nine out-of-distribution tasks in Table 6. The Optimism variant conducts optimistic planning instead of pessimistic planning in equation (9), which uses uniform sampling in the planner module. The Linear variant assumes a full connection to the causal matrix $M$, then directly uses linear MDP to parameterize the dynamics model $T$, which removes the causal discovery module in BECAUSE. The Full variant learns from the full batch of data to estimate the causal mask without iterative update. We report the results of the task-wise ablation with confidence interval and significance in Appendix Table 5 and 6.
@@ 5 Related Works
The objective mismatch in MBRL [43, 44] refers to the fact that pure MLE estimation of the world model does not align well with the control objective. Previous works [29, 45] propose reweighting during model training to alleviate this mismatch,  [46] proposes a goal-aware prediction by redistributing model error according to their task relevance. These works essentially reweight loss for the entire model training, while our work conducts reweighting just over the estimated causal mask more efficiently. More recently,  [9, 10] proposed a joint training between the world model and policies. Although joint optimization improves performance, they do not address the generalizability of the learned model under the distribution shift setting. In the offline setting, Model-based RL [2, 3, 4] employs model ensemble, pessimistic policy optimization or value iteration [47, 48], and an energy-based model for planning [49] to quantify uncertainty and improve test performance. To the best of our knowledge, no previous work explored or modeled the impact of distribution shift on the objective mismatch problem in MBRL.
@@
Recently, many RL algorithms have incorporated causality to improve reasoning capability [65] and generalizability. For instance, [66] and [67] explicitly estimate causal structures with the interventional data obtained from the environment in an online setting. These structures can be used to constrain the output space [19] or to adjust the buffer priority [68]. Building dynamic models in model-based RL [24, 69, 70] based on causal graphs is widely studied. Most existing causal MBRL works focus on estimating the causal world model by predicting transition dynamics and rewards. Existing methods learn this causal world model via sparsity regularization [23, 71], conditional independence test [24, 69, 72], variational inference [73, 12], counterfactual data augmentation [74, 75], hierarchical skill abstraction [76, 77], uncertainty quantification [40], reward redistribution [24, 78], causal context modeling [79, 80] and structure-aware state abstraction [12, 13, 81, 82] based on the controllability and task or reward relevance. However, the presence of confounders during data collection can skew the learned policy, making it susceptible to spurious correlations. Deconfounding solutions have been proposed either between actions and states [39, 83, 84] or among different dimensions of state variables [19, 85].
@@ A.1 Notation Summary
A.2 Derivation of Definition 3
@@ B.3 Proof of Lemma 3
Here, $I_{dd^{\prime}}$ represent the identity matrix of size $dd^{\prime}\times dd^{\prime}$, the last inequality holds based on the fact that $\|X\|_{2}\leq 1$ defined in equation (24). Since $\eta\leq\frac{1}{L}=1$, as a result, $\|I_{dd^{\prime}}-\eta X^{T}X\|_{2}=\max(1-\eta\lambda(X^{T}X))\leq 1$. We perform IHT for sufficient $K>0$ iterations and output the last step estimation as our solutions for either the offline dataset ${\mathcal{D}}$ (we use for practical optimization) or the perfect dataset ${\mathcal{D}}^{*}$. In practice, for any dataset such as ${\mathcal{D}}$ and any accuracy level $0\leq\xi\leq 1$, we have the output of IHT gradually converges to the optimal solution $\widehat{\beta}_{{\mathcal{D}}}^{M}$ of the problem in equation (LABEL:eq:lasso_final). Namely, after at most $K\simeq\log\left(\frac{\|M\|_{0}}{\xi}\right)$ steps [87, Corollary 1], the output satisfies $\left\|\widehat{\beta}_{{\mathcal{D}}}^{M}-\widehat{\beta}_{{\mathcal{D}}}^{M}(K)\right\|_{2}\leq\frac{\xi}{4}$. Similarly, based on the perfect dataset ${\mathcal{D}}^{*}$, the output of IHT gradually converges to $\widehat{\beta}_{{\mathcal{D}}^{*}}^{M}$ at the same rate. Consequently, the term of interest (a) can be bounded recursively as:
@@
The above proposition can be directly proved by applying [89, Lemma 17] over all $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$:
@@ C.4 Additional Experiment Results
We report the results of the task-wise performance of all baselines in the main experiment and variants in the ablation studies in Table 345, and  6.
@@ C.5 Additional Environment Description
Figure 8: Three environments used in this paper.

### v2

@@ Abstract
Offline model-based reinforcement learning (MBRL) enhances data efficiency by utilizing pre-collected datasets to learn models and policies, especially in scenarios where exploration is costly or infeasible. Nevertheless, its performance often suffers from the objective mismatch between model and policy learning, resulting in inferior performance despite accurate model predictions. This paper first identifies the primary source of this mismatch comes from the underlying confounders present in offline data for MBRL. Subsequently, we introduce BilinEar CAUSal rEpresentation (BECAUSE), an algorithm to capture causal representation for both states and actions to reduce the influence of the distribution shift, thus mitigating the objective mismatch problem. Comprehensive evaluations on 18 tasks that vary in data quality and environment context demonstrate the superior performance of BECAUSE over existing offline RL algorithms. We show the generalizability and robustness of BECAUSE under fewer samples or larger numbers of confounders. Additionally, we offer theoretical analysis of BECAUSE to prove its error bound and sample efficiency when integrating causal representation into offline MBRL. See more details in our project page: https://sites.google.com/view/be-cause.
@@ 1 Introduction
Offline Reinforcement Learning (RL) has shown great promise in learning directly from pre-collected datasets, especially in scenarios where active interaction is expensive or infeasible [1]. Specifically, offline model-based reinforcement learning (MBRL) [2, 3, 4], learning policies with an estimated world model, generally perform better than their model-free counterparts in long-horizon tasks such as self-driving vehicles [5], robotics [6], and healthcare [7]. However, offline RL suffers from distribution shift because the rollout data is either sampled from some suboptimal behavior policies or sampled from slightly different training environments compared to the deployment time [8].
@@
Although identifying distribution shift issues, many of the current offline MBRL works fail to model the shift in environment dynamics, which is ubiquitous and could cause catastrophic failure of trained policy at a slightly different deployment stage. Furthermore, since the learning objectives of the world models and policies are isolated from each other, a significant challenge in offline MBRL is objective mismatch [9, 10] problem (shown in Figure 1): models that achieve a lower training loss are not necessarily better for control performance. For example, a dynamics model achieve relatively low prediction loss, yet it may not be sufficient to guide the planner to the high-reward region. Previous works [9, 10] have attempted to reduce such objective mismatch by jointly learning the model and policy. However, the performance is suboptimal due to the lack of digging into the underlying cause of the mismatch issue [8].
In this work, we identify that the objective mismatch between model estimation and policy learning comes from two sources of distribution shift in offline MBRL: (1) shift between the online optimal policy and offline sub-optimal behavior policies, and (2) shift between the data collection environment and online testing environments. Unlike humans, who make decisions based on reasoning over task-relevant factors, models in offline RL memorize correlations without learning the causality. The sub-optimal behavior policies introduce spurious correlations [11] between actions and states, making the model memorize specific actions. When online testing environments differ from the data collection environment, the model could overfit spurious correlations in the state and fail to generalize to unseen states. Based on the analysis of the above mismatch, our work differs from previous work of causal model-based RL [12, 13] in that we model causality in both model and policy learning., We aim to avoid spurious correlation by discovering underlying structures between abstracted states and actions.
@@ 3.1 Causal Representation for ASC-MDP
The assumption 1 can also be interpreted as task independence in  [24], invariant state representation [25], or invariant action effect in [26]. See detailed comparison in appendix Table 4.
@@ 4 Experiment Results
Figure 4: Three environments used in this paper.
In this section, we conduct a comprehensive empirical evaluation of BECAUSE’s generalization performance in a diverse set of environments, covering different decision-making problems in the grid world, manipulation, and autonomous driving domains, shown in Figure 4.
@@ 4.1 Experiment Setting
We design 18 tasks in 3 representative RL environments in Figure 4. Agents need to acquire reasoning capabilities to receive higher rewards and achieve goals.
@@ 4.2 Experiment Results Analysis
  • RQ3: How does BECAUSE scale up to visual RL tasks with image observation input compared to other visual RL baselines?
  • RQ4: How does BECAUSE achieve the aforementioned generalizability by mitigating the objective mismatch problem in offline MBRL?
@@
Table 1: Comparison of visual RL performance.
Tasks ICIL IFactor BECAUSE
Unlock-I-random 0.8$\pm$0.8 4.3$\pm$1.1 15.7$\pm$3.3
Unlock-O-random 1.5$\pm$1.8 4.7$\pm$1.6 5.9$\pm$0.9
Unlock-I-medium 5.3$\pm$2.0 30.2$\pm$4.1 62.0$\pm$4.6
Unlock-O-medium 8.6$\pm$4.2 15.4$\pm$2.4 71.6$\pm$9.1
Unlock-I-expert 8.7$\pm$3.4 34.0$\pm$4.8 63.7$\pm$3.9
Unlock-O-expert 17.1$\pm$4.2 16.7$\pm$3.1 73.6$\pm$19.5
For RQ3, we conduct experiments with visual inputs in the Unlock environments with ICIL [38] and IFactor [13]. We parameterize the feature encoder as a three-layer CNN with 128 dimensions hidden size for all the baselines and our methods. The results in Table 1 show that BECAUSE can significantly improve both in-distribution and out-of-distribution performance under different quality of behavior policies.
@@
For RQ4, we aim to understand whether BECAUSE achieves higher performance by resolving the objective mismatch problem. We first collect two groups of trajectories: $\bm{\tau_{pos}}$ and $\bm{\tau_{neg}}$, each with positive reward (success) and negative reward (failure) in Unlock task with sparse goal-reaching reward. We want to have a model whose loss is informative for discriminating control results, that is, we wish ${\mathcal{L}}_{model}({\color[rgb]{0.44,0.72,0.93}\bm{\tau_{pos}}})<{\mathcal{L}}_{model}({\color[rgb]{0.96,0.49,0.43}\bm{\tau_{neg}}})$. According to our visualization in Figure 6, in Unlock-Expert and Unlock-Medium, the ratio of $\bm{\tau_{pos}}$ is much higher in BECAUSE than MOPO among the trajectories with low model loss. In Unlock-Random, the mismatch of the model and control objective is more significant, since the demonstration is poor in state coverage. MOPO cannot succeed even when the model loss is low, whereas our methods can. We perform a hypothesis test with $H_{0}:{\mathcal{L}}_{model}({\color[rgb]{0.44,0.72,0.93}\bm{\tau_{pos}}})<{\mathcal{L}}_{model}({\color[rgb]{0.96,0.49,0.43}\bm{\tau_{neg}}})$. In BECAUSE, this desired property is more significant than MOPO attributed to the causal representation we learn, indicating a reduction of objective mismatch. We attach detailed discussions for the mismatch evaluation in Appendix C.3 and Table 5.
4.3 Ablation Studies
Table 2: The ablation studies between BECAUSE and its variants. We report the overall Success rate (%) over 9 in-distribution (I) and 9 out-of-distribution (O) tasks, respectively. Bold is the best.
@@ 4.3 Ablation Studies
We conducted ablation studies with three variants of BECAUSE and report the average success rate across nine in-distribution and nine out-of-distribution tasks in Table 2. The Optimism variant conducts optimistic planning instead of pessimistic planning in Equation (9), which uses uniform sampling in the planner module. The Linear variant assumes a full connection to the causal matrix $M$, then directly uses linear MDP to parameterize the dynamics model $T$, which removes the causal discovery module in BECAUSE. The Full variant learns from the full batch of data to estimate the causal mask without iterative update. We report the results of the task-wise ablation with confidence interval and significance in Appendix Table 8 and 9.
@@ 5 Related Works
The objective mismatch in MBRL [43, 44] refers to the fact that pure MLE estimation of the world model does not align well with the control objective. Previous works [29, 45] propose reweighting during model training to alleviate this mismatch,  [46] proposes a goal-aware prediction by redistributing model error according to their task relevance. These works essentially reweight loss for the entire model training, while our work conducts reweighting just over the estimated causal mask more efficiently. More recently,  [9, 10] proposed a joint training between the world model and policies. Although joint optimization improves performance, they do not address the generalizability of the learned model under the distribution shift setting. In the offline setting, Model-based RL [2, 3, 4, 47] employs model ensemble, pessimistic policy optimization or value iteration [48, 49], and an energy-based model for planning [50] to quantify uncertainty and improve test performance. To the best of our knowledge, no previous work explored or modeled the impact of distribution shift on the objective mismatch problem in MBRL.
@@
Recently, many RL algorithms have incorporated causality to improve reasoning capability [66] and generalizability. For instance, [67] and [68] explicitly estimate causal structures with the interventional data obtained from the environment in an online setting. These structures can be used to constrain the output space [19] or to adjust the buffer priority [69]. Building dynamic models in model-based RL [24, 70, 71] based on causal graphs is widely studied. Most existing causal MBRL works focus on estimating the causal world model by predicting transition dynamics and rewards. Existing methods learn this causal world model via structural regularization [23, 72, 73], conditional independence test [24, 70, 74], variational inference [75, 12], counterfactual data augmentation [76, 77], hierarchical skill abstraction [78, 79], uncertainty quantification [40], reward redistribution [24, 80], causal context modeling [81, 82] and structure-aware state abstraction [12, 13, 83, 84] based on the controllability and task or reward relevance. However, the presence of confounders during data collection can skew the learned policy, making it susceptible to spurious correlations. Deconfounding solutions have been proposed either between actions and states [39, 85, 86] or among different dimensions of state variables [19, 87].
@@ 6 Conclusion
Acknowledgement
This work is partially supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR00112320012. The work of L. Shi is supported in part by the Resnick Institute and Computing, Data, and Society Postdoctoral Fellowship at California Institute of Technology.
@@ A.1 Notation Summary
A.2 Equivalence of the Assumptions with Previous Works
Table 4: Summary of assumptions and equivalent forms.
Assumption Base and Equivalent Format
Invariant State Abstraction [25] $T(s^{\prime}|s,a;u)=T(\phi(s^{\prime})|\phi(s),a)\cdot p_{e}(u^{\prime}|s,u)$
$T(s^{\prime}|s,a;u)\propto T(\phi(s^{\prime})|\phi(s),a)$
Task Independence [24] $T(\phi(s^{\prime})|\phi(s),a)=T(s^{\prime}_{C}|s_{C},s_{R},a)p(s^{\prime}_{R}|s_{R})$
$T(\phi(s^{(1)})|\phi(s^{(1)}),a;u^{(1)})=T(\phi(s^{(2)})|\phi(s^{(2)}),a;u^{(2)})$
Invariant Action Effect [26] $T_{\text{online}}(s^{\prime}|s,a)=\sum_{u}T(s^{\prime}|s,a,u)\hat{p}(u|s)$
$T_{\text{offline}}(s^{\prime}|s,a)=\sum_{u}T(s^{\prime}|s,a,u)\hat{p}(u|s)$
$T_{\text{online}}(s^{\prime}|s,a)\propto T_{\text{offline}}(s^{\prime}|s,a)$
Invariant Causal Graph ${u}^{(1)}\neq{u}^{(2)},{M}(\mathbf{u}^{(1)})={M}({u}^{(2)})$
$T(s^{\prime}|s,a;u^{(1)})=T(s^{\prime}|s,a;u^{(2)})$
A.3 Derivation of Definition 3
@@ B.3 Proof of Lemma 3
Here, $I_{dd^{\prime}}$ represent the identity matrix of size $dd^{\prime}\times dd^{\prime}$, the last inequality holds based on the fact that $\|X\|_{2}\leq 1$ defined in equation (24). Since $\eta\leq\frac{1}{L}=1$, as a result, $\|I_{dd^{\prime}}-\eta X^{T}X\|_{2}=\max(1-\eta\lambda(X^{T}X))\leq 1$. We perform IHT for sufficient $K>0$ iterations and output the last step estimation as our solutions for either the offline dataset ${\mathcal{D}}$ (we use for practical optimization) or the perfect dataset ${\mathcal{D}}^{*}$. In practice, for any dataset such as ${\mathcal{D}}$ and any accuracy level $0\leq\xi\leq 1$, we have the output of IHT gradually converges to the optimal solution $\widehat{\beta}_{{\mathcal{D}}}^{M}$ of the problem in equation (LABEL:eq:lasso_final). Namely, after at most $K\simeq\log\left(\frac{\|M\|_{0}}{\xi}\right)$ steps [89, Corollary 1], the output satisfies $\left\|\widehat{\beta}_{{\mathcal{D}}}^{M}-\widehat{\beta}_{{\mathcal{D}}}^{M}(K)\right\|_{2}\leq\frac{\xi}{4}$. Similarly, based on the perfect dataset ${\mathcal{D}}^{*}$, the output of IHT gradually converges to $\widehat{\beta}_{{\mathcal{D}}^{*}}^{M}$ at the same rate. Consequently, the term of interest (a) can be bounded recursively as:
@@
The above proposition can be directly proved by applying [91, Lemma 17] over all $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$:
@@ C.4 Additional Experiment Results
We report the results of the task-wise performance of all baselines in the main experiment and variants in the ablation studies in Table 678, and  9.